OpenStudy (anonymous):

-9e^(7x) over 7x-3 find derivative!

OpenStudy (anonymous):

\[\frac{-9e^{7x}}{7x-3}\]

OpenStudy (anonymous):

use quotient rule and chain rule. quotient rule is \[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\] with \[f(x)=e^{7x},f'(x)=7e^{7x},g(x)=7x-3,g'(x)=7\] and you can multiply by that annoying -9 at the end

OpenStudy (anonymous):

you get \[-9\times\frac{(7x-3)\times 7e^{7x}-e^{7x}\times 7}{(7x-3)^2}\]

OpenStudy (anonymous):

you didnt finish the other one

OpenStudy (anonymous):

what other one?

OpenStudy (anonymous):

the problem that I posted

OpenStudy (anonymous):

i can't find it. link

OpenStudy (anonymous):

i already solved it but apparently is wrong

OpenStudy (anonymous):

i wrote it out